# ROODI: Reconstructing Occluded Objects with Denoising Inpainters

### v1

@@ Abstract
While the quality of novel-view images has improved dramatically with 3D Gaussian Splatting, extracting specific objects from scenes remains challenging. Isolating individual 3D Gaussian primitives for each object and handling occlusions in scenes remain far from being solved. We propose a novel object extraction method based on two key principles: (1) being object-centric by pruning irrelevant primitives; and (2) leveraging generative inpainting to compensate for missing observations caused by occlusions. For pruning, we analyze the local structure of primitives using K-nearest neighbors, and retain only relevant ones. For inpainting, we employ an off-the-shelf diffusion-based inpainter combined with occlusion reasoning, utilizing the 3D representation of the entire scene. Our findings highlight the crucial synergy between pruning and inpainting, both of which significantly enhance extraction performance. We evaluate our method on a standard real-world dataset and introduce a synthetic dataset for quantitative analysis. Our approach outperforms the state-of-the-art, demonstrating its effectiveness in object extraction from complex scenes. Project page is available at https://yeonjin-chang.github.io/ROODI/
@@
1 Introduction
The quality of novel view synthesis from a set of captured images has improved dramatically in recent years, largely due to the advent of Neural Radiance Fields (NeRF) [29], which was later superseded by 3D Gaussian Splatting (3DGS) [18] for its superior rendering efficiency. These methods represent a scene as a radiance field, encoded either as a neural network [29], a hash grid [32], or a set of 3D Gaussians [18], and render it via volume rendering [16] or splatting [67]. Numerous extensions have been introduced to enhance quality [1, 48, 62], support dynamic scenes [36, 27], enable 3D content generation [34, 45], and incorporate semantic understanding [61, 3]. Among the many applications of 3DGS, we focus on those aimed at extracting specific objects from a scene [57, 9, 3, 4, 66, 59].
Among the many, Gaussian Grouping [57] focuses on distilling 2D semantic segmentation maps into semantic features for 3D Gaussians, which are later classified to form groups of 3D Gaussians. OmniSeg3D [59], SAGA [3], and ClickGaussian [9] also leverage 2D semantic maps, but instead use them as masks to train the semantic features of 3D Gaussians via contrastive learning, which is subsequently used for object extraction. LBG [4] explores integrating features from image foundation models into 3D Gaussians by applying 2D semantic maps as masks. However, as shown in Fig. 1, none of the existing methods yield a clearly extracted target model and have so-called floaters due to other nearby objects, and suffer from occlusions.
@@ 1 Introduction
In this work, we propose a novel method to extract specific objects from a 3DGS model, based on two key ideas: (1) objects are local—we are interested in a single cluster of primitives that form a single object, and (2) occlusions can easily be reasoned by comparing the primitives of the target object with those of the full scene. Using the former, we look into the local structure of the primitives while reconstructing the target object, and prune those that are not part of the target object—we implement this by looking at the distance to the K-th nearest neighbor of each primitive. Using the latter, we use a pretrained diffusion model to inpaint the occluded regions of the target object. As shown in Fig. 1, this leads to clear object extraction, one that is robust to occlusions and floaters.
@@
In more detail, we start with a 3DGS model of the scene. We use Segment Anything Model 2 (SAM2) [23, 37] to obtain the segmentation map of the object of interest. As in Gaussian Grouping [57], we then distill the segmentation map into the 3DGS model through a brief training period of 500 iterations. We then update 3D Gaussians specific to the target object by: (1) pruning Gaussians whose K-th nearest neighbor exceeds a certain distance (hence floaters), (2) rendering both the target object and the full scene to identify occluded regions of the target object, and (3) inpainting the occluded regions of the target object with a pretrained inpainting diffusion model [33, 46], and (4) updating the parameters of the 3D Gaussians to better match the inpainted target object. Notably, all of these steps are crucial, as without pruning floaters, the diffusion model is unable to recognize or inpaint the target object due to floaters, and inpainting is further critical to obtain clear, sharp images—we show empirically that simple general diffusion-based enhancement is insufficient.
@@
To evaluate our method, we conduct experiments on both synthetic and real-world datasets. We use the standard LERF dataset [19] to allow qualitative comparison with existing methods. As the real-world dataset does not have a ground-truth object-centric extraction (e.g., parts of object can be occluded), we create a new synthetic dataset, named MultiObjectBlender. We render images from a 3D scene composed of many objects from ShapeNet [5] using Blender [2], where many objects are occluding each other. We then render both the full scene and only the target object, the latter of which can be used to quantitatively evaluate the extraction quality. Our method significantly outperforms existing methods on both datasets.
Contributions.
To summarize, in this work:
@@
  • we propose a method to extract objects within a 3DGS model that prunes floaters through the K-th nearest neighbor, and performs occlusion reasoning and inpainting;
  • we demonstrate both the removal of floaters and the inpainting of occluded regions are important;
@@ 2 Related Work
Research on 3D scene reconstruction from multi-view images has gained momentum since NeRF [29], which introduced an implicit representation of 3D space using neural networks, enabling photo-realistic novel view synthesis. However, NeRF suffers from long training and inference times, which led to subsequent research incorporating voxel grids [60, 10, 14, 44], tri-planes [20, 11, 6], and hybrid representations [47, 65].
@@
In contrast to the implicit formulation of NeRFs, 3D Gaussian Splatting (3DGS) [18] emerged as an explicit representation that models the 3D radiance field using Gaussian primitives. This approach significantly reduces rendering time and enables real-time rendering while achieving high visual quality. Following the introduction of 3DGS, research has rapidly expanded across various topics and applications, including dynamic scenes with moving objects [27, 25, 55, 56, 52], 3D/4D object generation [45, 58, 8, 38, 63, 54, 12], and semantic encoding in representations [3, 57, 15, 7]. Among these, extracting specific objects within a 3D scene using semantic encoding has become an important topic for scene editing. Existing methods, however, often struggle with floating artifacts or occlusions as shown in Fig. 1, making it challenging to achieve clean and accurate extractions. In this paper, we address the problem of extracting objects from 3D space while preserving high-quality details.
@@
Figure 2: Overview – (a) Given a standard 3DGS model along with multi-view images, we select the object to extract. (b) We obtain segmentation maps for the target objects using SAM2, (c) to distill the semantic map into the object feature of 3D Gaussians. (d) Using the object features, we initially separate the object and prune floaters, thresholding by the distance to K-th nearest Gaussian. (e) We then utilize an inpainting diffusion model to fill in the missing parts, reasoning occlusions based on the depth map.
3D semantic segmentation.
@@
We first briefly review semantic segmentation using 3DGS, as many object extraction methods rely on it. Segmentation on 3DGS is typically achieved by incorporating an additional semantic feature field [9, 66, 57, 3, 41], or assigning 3D labels [40, 4]. These semantic features are trained using 2D semantic maps obtained from foundation models [23, 24] for 2D segmentation. Some distill 2D semantic features to 3D semantic features through contrastive learning [9, 3, 59, 41], while others directly train 3D features using a multichannel renderer [66]. Alternatively, other methods [4, 40] map 2D semantic masks to 3D labels, and Ye et al. [57] directly trains semantic features by rendering them and performing pixel-wise classification under the supervision of 2D semantic maps. These methods achieve strong semantic segmentation performance when evaluating their rendered 2D semantic maps against those obtained from 2D foundation models. However, they place less emphasis on segmentation in 3D space itself, which is directly related to extraction quality.
3D object extraction.
@@
Many 3D segmentation studies present object extraction as a downstream application, and not as a primary research focus [57, 9, 4, 59, 22, 3, 41]. Since these methods include semantic features, they can extract objects either by using a classifier to distinguish the groups to which primitives belong [57] or by applying thresholding based on feature similarity to extract primitives with high similarity [3, 59]. However, most approaches do not output an actual extracted model but select primitives with similar features, often leaving the threshold for the similarity undefined. This can lead to suboptimal extraction outcomes when used as an object extractor, and can perform poorly when the scene is complex with occlusions. In this paper, we prioritize achieving high-quality extracted models with inpainted occluded regions.
Inpainting 3D scenes from 2D projections.
Inpainting a 3D scene is often reformulated as inpainting the multi-view 2D projections of the existing 3D content [51, 30, 35, 49, 31]. State-of-the-art 2D inpainting methods are typically based on diffusion models [42, 43, 13], with some work directly using pretrained image diffusion models [26, 17]. They iteratively add noise to the whole image and stitch the denoised area-to-inpaint with the preserved area so that the image is inpainted in a harmonizing way. Other methods modify and finetune pretrained image diffusion models to inpaint an image [39, 33]. In this work, we use an off-the-shelf finetuned inpainting diffusion model, “SD-XL Inpainting 0.1”  [46].
@@ 3 Method
We provide an overview of our method in Fig. 2. To start with, we pretrain a 3DGS model of the entire scene. Then, for the objects of interest, single or multiple, we use Segment Anything V2 (SAM2) [37] with just one to two clicks to obtain their segmentation maps from multiple views. With these masks, we find the subset of Gaussians that correspond to the target object(s) by distilling these masks into the pretrained 3D Gaussians as in Gaussian Grouping [57]. We then look at the distilled object features to obtain object-specific 3D Gaussians. These Gaussians, however, are crude and contain many artifacts and holes as shown in Fig. 2d. The core of our contribution is how we further train these Gaussians to properly represent objects as shown in Fig. 2e. We first discuss the initial training process of 3DGS and how we distill masks into the Gaussians in Sec. 3.1. We then detail our core contributions, object-centric pruning (Sec. 3.2) and how we utilize pre-trained inpainting models with occlusion reasoning (Sec. 3.3).
@@ 3.1 Initial training
We begin by following the standard training process of 3D Gaussian Splatting (3DGS) [18] to construct a 3DGS model of the entire scene, with the more-recent Markov Chain Monte Carlos-based solver [21]. We parameterize each Gaussian by a rotation matrix $R\in^{3\times 3}$, scaling $S\in$, position $X\in^{3}$, opacity $\alpha\in$, and color feature $F_{c}\in^{h}$, where $h$ is the number of spherical harmonic coefficients. We train using a set of images =$\mathcal{I}=\{I^{1},I^{2},...,I^{t}\}$, where $I^{*}\in^{H\times W}$. To render a pixel $p\in^{2}$, we first sort Gaussians by their depth relative to the camera (rendering view), and then composite using $\alpha$-blending. We compute the color of pixel $p$ as:
@@
where $N\in$ is the number of Gaussians contributing to the pixel’s color, $c_{i}\in^{3}$ and $\tilde{\alpha}_{i}\in$ represent the color and the opacity of the $i$-th Gaussian at pixel $p$ after being projected into the rendering view.
The parameters of the Gaussians are updated to minimize both the $\mathcal{L}_{1}$ loss and the D-SSIM loss [18], where the former is a per-pixel loss enforcing colors to be the same, and the latter focuses on the local structure:
@@
We then, as in Gaussian Grouping [57] and many others that incorporate semantics to Gaussians, distill the masks as an additional feature to each Gaussian. Specifically, for each Gaussian, we add an object feature $F_{o}\in^{d}$ that encodes the object information of the $C$ objects, where $d$ denotes the dimension of the object features. Then, after rendering the features of the Gaussians, we pass these features through a shallow classifier to generate the object feature map $\hat{O}\in^{H\times W\times C}$ of size $H\times W$, as shown in Fig. 2c. We then compare this feature map with the segmentation masks obtained from SAM2 [37], and minimize the cross entropy between the object feature map $\hat{O}$ and the segmentation masks $O\in^{H\times W}$ that contain which class each pixel belongs to:
$\mathcal{L}_{\text{obj}}=-\frac{1}{HW}\sum_{i=1}^{H}\sum_{j=1}^{W}\sum_{c=1}^{C}\mathbbm{1}(O_{i,j}=c)\log\hat{O}_{c,i,j},$ (3)
where $\mathbbm{1}$ is the indicator function. After 500 iterations of training, the object feature becomes sufficiently trained, making it possible to reasonably determine which Gaussians belong to which objects.
@@ 3.2 Object-centric pruning
A key observation in reconstructing a specific object is that the task itself is about a single object. However, directly distilling masks or semantics often leads to the distillation being spread across multiple Gaussians. This is because the Gaussians are typically non-opaque, meaning multiple Gaussians must work together to form surfaces. This, in turn, results in the appearance of floaters, as shown in Fig. 3. Given that object-specific 3D Gaussians should represent only the target object, we can effectively ‘prune’ these floaters by simply removing those that are not structurally connected.
Rendering Inpaint mask Inpainted
w/o Pruning
Refer to caption Refer to caption Refer to caption
Our Method
Refer to caption Refer to caption Refer to caption
Figure 3: Pruning affects inpainting quality – Gaussian segmentation often misclassifies Gaussians, and these appear as floaters. These floaters can, in fact, cause inpainters to do harm rather than help. In this example, inpainting results in an extra eye (denoted as red arrow), with the floaters being understood as shadows. With our pruning, this can be avoided.
Specifically, for each Gaussian, we look at its distance to its K-th nearest Gaussian, and prune it if the distance is larger than a threshold $\epsilon$. Mathematically, we prune those that satisfy
$\|X_{i}-X_{K_{i}}\|_{2}>\epsilon,$ (4)
where $X_{i}$ is the position of the $i$-th Gaussian, and $X_{K_{i}}$ is the position of the $K$-th nearest Gaussian for the $i$-th Gaussian. In other words, if a Gaussian, considering $\epsilon$ as the connectivity threshold, is not connected to at least $K$ other Gaussians, we prune it. This process can be effectively implemented as a CUDA kernel, allowing the process to be completed within 20 seconds for each scene in our test set. We run pruning once after 500 iterations, before inpainting starts. This results in an average of 1.5% of the Gaussians being pruned.
We show in Fig. 3 how our pruning removes floaters. What is more interesting is that these floaters heavily impact the quality of what off-the-shelf inpainting models can do, as floaters are not what these models have been trained to deal with. But with our pruning, and with the floaters removed, inpainting results are improved as shown, which ultimately contributes to better final object reconstruction.

### v2

@@ Abstract
While the quality of novel-view images has improved dramatically with 3D Gaussian Splatting, extracting specific objects from scenes remains challenging. Isolating individual 3D Gaussian primitives for each object and handling occlusions in scenes remains far from being solved. We propose a novel object extraction method based on two key principles: (1) object-centric reconstruction through removal of irrelevant primitives; and (2) leveraging generative inpainting to compensate for missing observations caused by occlusions. For pruning, we propose to remove irrelevant Gaussians by looking into how close they are to its K-nearest neighbors and removing those that are statistical outliers. Importantly, these distances must take into account the actual spatial extent they cover—we thus propose to use Wasserstein distances. For inpainting, we employ an off-the-shelf diffusion-based inpainter combined with occlusion reasoning, utilizing the 3D representation of the entire scene. Our findings highlight the crucial synergy between proper pruning and inpainting, both of which significantly enhance extraction performance. We evaluate our method on a standard real-world dataset and introduce a synthetic dataset for quantitative analysis. Our approach outperforms the state-of-the-art, demonstrating its effectiveness in object extraction from complex scenes.
1 Introduction
3D scenes are typically composed of numerous objects. Naturally, beyond the original 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023), several works attempt to interpret 3D scenes at the object level to better capture their semantic structure (Ye et al. 2024; Choi et al. 2024; Cen et al. 2025; Chacko et al. 2025; Zhou et al. 2024; Ying et al. 2024). These studies primarily train semantic features of 3D scenes by leveraging semantic maps from 2D segmentation models, modeling the overall structure and semantics of the scene. However, extracting the 3DGS model of target objects with these semantic features introduces additional challenges—precise separation of the target object from the rest of the scene is difficult and (self-)occlusions further make this problem harder. As shown in Fig. 1, none of the existing methods yield a clearly extracted target model and have so-called floaters.
@@ 1 Introduction
In this work, we propose a novel method for 3D object extraction, based on three key ideas: (1) regardless of the choice of which object, scene geometry must remain the same—the full scene geometry serves as base knowledge for object extraction, (2) objects are local—we are interested in a single cluster of primitives that form a single object, and (3) occlusions can easily be reasoned by comparing the primitives of the target object with those of the full scene. Based on (1), we adopt a two-stage pipeline that trains the scene geometry first, followed by additional training of object features to distinguish objects. Based on (2), we look into the local structure of the primitives and prune those that are not part of the target object. We propose to use the distance of each Gaussian to their nearest neighbors and remove those that are statistical outliers. As Gaussians are entities that cover a region in 3D space, we consider the distances in terms of Wasserstein distances—we show later in Sec. 4.3 that this is important for proper pruning. From (3), we use a pretrained diffusion model to inpaint the occluded regions of the target object. This leads to clear object extraction, one that is robust to occlusions and floaters; see Fig. 1.
@@
In more detail, we first train a 3DGS model of the scene, then extract the 3DGS model of the target objects in the second stage. To extract an object, we start with Segment Anything Model 2 (SAM2) (Kirillov et al. 2023; Ravi et al. 2024) to obtain the segmentation map of the object of interest. We then distill this map into the 3DGS model by briefly fine-tuning the model for 500 iterations. We then update 3D Gaussians specific to the target object by: (1) pruning non-object Gaussians by removing those that are statistical outliers when looking into their Wasserstein distances to nearest neighbors; (2) rendering both the target object and the full scene to identify occluded regions of the target object; and (3) inpainting the occluded regions of the target object with a pretrained inpainting diffusion model (Podell et al. 2023; The Diffusers team 2023); and finally (4) updating the parameters of the 3D Gaussians to better match the inpainted target object. Notably, all of these steps are crucial, as without pruning floaters, the diffusion model is unable to recognize or inpaint the target object due to floaters, and inpainting is further critical to obtain clear, sharp images—we show empirically that simple general diffusion-based enhancement is insufficient.
@@
To evaluate our method, we conduct experiments on both synthetic and real-world datasets. We use the standard LERF dataset (Kerr et al. 2023) to allow qualitative comparison with existing methods. As the real-world dataset does not have a ground-truth object-centric extraction (e.g., parts of object can be occluded), we create a new synthetic dataset, named MultiObjectBlender. We render images from a 3D scene composed of many objects from ShapeNet (Chang et al. 2015) using Blender (Blender Online Community 2024), where many objects are occluding each other. We then render both the full scene and only the target object, the latter of which can be used to quantitatively evaluate the extraction quality. Our method significantly outperforms existing methods on both datasets.
Contributions. To summarize, in this work:
@@
  • we propose a method to extract objects within a 3DGS model that prunes non-object Gaussians through statistical outlier removal, and performs occlusion reasoning and inpainting;
  • we propose to use Wasserstein distances when measuring distances between Gaussians for this purpose;
  • we demonstrate that the two-stage pipeline, the removal of non-object Gaussians, and the inpainting of occluded regions are all essential;
@@ 2 Related Work
3D reconstruction. Research on 3D scene reconstruction from multi-view images has gained momentum since NeRF (Mildenhall et al. 2020), which introduced an implicit representation of 3D space using neural networks, enabling photo-realistic novel view synthesis. However, NeRF suffers from long training and inference times, which led to subsequent research incorporating voxel grids (Yu et al. 2021; Fridovich-Keil et al. 2022; Hu et al. 2022; Sun, Sun, and Chen 2022), tri-planes (Khatib and Giryes 2024; Fridovich-Keil et al. 2023; Chen et al. 2022), and hybrid representations (Turki et al. 2024; Zhang, Chen, and Cui 2025).
@@
In contrast to the implicit formulation of NeRFs, 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) emerged as an explicit representation that models the 3D radiance field using Gaussian primitives. This approach enables real-time rendering while achieving high visual quality. Following the introduction of 3DGS, research has rapidly expanded across various topics and applications, including dynamic scenes with moving objects (Luiten et al. 2024; Li et al. 2024; Yang et al. 2023, 2024; Wu et al. 2024a), 3D/4D object generation (Tang et al. 2023; Yi et al. 2024; Chen et al. 2024b; Ren et al. 2023; Zeng et al. 2024; Wu et al. 2024c; Gao et al. 2024), and semantic encoding in representations (Cen et al. 2025; Ye et al. 2024; Ji et al. 2024; Chen et al. 2024a). Among these, extracting specific objects within a 3D scene using semantic encoding has become an important topic for scene editing. Existing methods, however, often struggle with floating artifacts or occlusions as shown in Fig. 1, making it challenging to achieve clean and accurate extractions. In this paper, we address the problem of extracting objects from 3D space while preserving high-quality details.
@@
Figure 2: Overview – (a) Given a standard 3DGS model along with multi-view images, we select the object to extract. (b) We obtain segmentation maps for the target objects using SAM2, (c) to distill the semantic map into the object feature of 3D Gaussians. (d) Using the object features, we initially separate the object and prune floaters, via our statistical pruning with Wasserstein distance. (e) We then utilize an inpainting diffusion model to fill in the missing parts, reasoning occlusions based on the depth map.
@@
3D semantic segmentation. We first briefly review semantic segmentation using 3DGS, as many object extraction methods rely on it. Segmentation on 3DGS is typically achieved by incorporating an additional semantic feature field (Choi et al. 2024; Zhou et al. 2024; Ye et al. 2024; Cen et al. 2025; Silva et al. 2024), or assigning 3D labels (Shen, Yang, and Wang 2024; Chacko et al. 2025). These semantic features are trained using 2D semantic maps obtained from foundation models (Kirillov et al. 2023; Li et al. 2022) for 2D segmentation. Some distill 2D semantic features to 3D semantic features through contrastive learning (Choi et al. 2024; Cen et al. 2025; Ying et al. 2024; Silva et al. 2024), while others directly train 3D features using a multichannel renderer (Zhou et al. 2024). Alternatively, other methods (Chacko et al. 2025; Shen, Yang, and Wang 2024) map 2D semantic masks to 3D labels, and Ye et al. (2024) directly trains semantic features by rendering them and performing pixel-wise classification under the supervision of 2D semantic maps. These methods achieve strong semantic segmentation performance when evaluating their rendered 2D semantic maps against those obtained from 2D foundation models. However, they place less emphasis on segmentation in 3D space itself, which is directly related to extraction quality.
@@
3D object extraction. Many 3D segmentation studies present object extraction as a downstream application, and not as a primary research focus (Ye et al. 2024; Choi et al. 2024; Chacko et al. 2025; Ying et al. 2024; Kim et al. 2024; Cen et al. 2025; Silva et al. 2024). Since these methods include semantic features, they can extract objects either by using a classifier to distinguish the groups to which primitives belong (Ye et al. 2024) or by applying thresholding based on feature similarity to extract primitives with high similarity (Cen et al. 2025; Ying et al. 2024). However, most approaches do not output an actual extracted model but select primitives with similar features, often leaving the threshold for the similarity undefined. This can lead to suboptimal extraction outcomes when used as an object extractor, and can perform poorly when the scene is complex with occlusions. In this paper, we prioritize achieving high-quality extracted models with inpainted occluded regions.
@@ 3 Method
We provide an overview of our method in Fig. 2. To start with, we pretrain a 3DGS model of the entire scene. Then, for the objects of interest, single or multiple, we use Segment Anything V2 (SAM2) (Ravi et al. 2024) with just one to two clicks to obtain their segmentation maps from multiple views. With these masks, we find the subset of Gaussians that correspond to the target object(s) by distilling these masks into the object features of pretrained 3D Gaussians. We then look at the distilled object features to obtain object-specific 3D Gaussians. These Gaussians, however, are crude and contain many artifacts and holes as shown in Fig. 2d. The core of our contribution is how we further train these Gaussians to properly represent objects as shown in Fig. 2e.
This section is organized as follows: we first discuss the initial training process of 3DGS and how we distill masks into the Gaussians in Sec. 3.1; we then detail our core contributions, object-centric pruning (Sec. 3.2) and how we utilize pre-trained inpainting models with occlusion reasoning (Sec. 3.3).
@@ 3.1 Initial training
We begin by following the standard training process of 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) to construct a 3DGS model of the entire scene, with the more-recent Markov Chain Monte Carlos-based solver (Kheradmand et al. 2025). We parameterize each Gaussian by a rotation matrix $R\in\mathds{R}^{3\times 3}$, scaling $S\in\mathds{R}$, position $X\in\mathds{R}^{3}$, opacity $\alpha\in\mathds{R}$, and color feature $F_{c}\in\mathds{R}^{h}$, where $h$ is the number of spherical harmonic coefficients. We train using a set of images $\mathcal{I}=\{I^{1},I^{2},...,I^{t}\}$, where $I^{*}\in\mathds{R}^{H\times W}$. To render a pixel $p\in\mathds{R}^{2}$, we first sort Gaussians by their depth relative to the camera (rendering view), and then composite using $\alpha$-blending. We compute the color of pixel $p$ as:
@@
where $N\in\mathds{R}$ is the number of Gaussians contributing to the pixel’s color, $c_{i}\in\mathds{R}^{3}$ and $\tilde{\alpha}_{i}\in\mathds{R}$ represent the color and the opacity of the $i$-th Gaussian at pixel $p$ after being projected into the rendering view.
The parameters of the Gaussians are updated to minimize both the $\mathcal{L}_{1}$ loss and the D-SSIM loss (Kerbl et al. 2023), where the former is a per-pixel loss enforcing colors to be the same, and the latter focuses on the local structure:
@@
We then, as in Gaussian Grouping (Ye et al. 2024) and many others that incorporate semantics to Gaussians, distill the masks as an additional feature to each Gaussian. Specifically, for each Gaussian, we add an object feature $F_{o}\in\mathds{R}^{d}$ that encodes the object information of the $C$ objects, where $d$ denotes the dimension of the object features. Then, after rendering the features of the Gaussians, we pass these features through a shallow classifier to generate the object feature map $\hat{O}\in\mathds{R}^{H\times W\times C}$ of size $H\times W$, as shown in Fig. 2c. We then compare this feature map with the segmentation masks obtained from SAM2 (Ravi et al. 2024), and minimize the cross entropy between the object feature map $\hat{O}$ and the segmentation masks $O\in\mathds{R}^{H\times W}$ that contain which class each pixel belongs to:
$\mathcal{L}_{\text{obj}}=-\frac{1}{HW}\sum_{i=1}^{H}\sum_{j=1}^{W}\sum_{c=1}^{C}\mathds{1}(O_{i,j}=c)\log\hat{O}_{c,i,j},$ (3)
where $\mathds{1}$ is the indicator function. After 500 iterations of training, the object feature becomes sufficiently trained, making it possible to reasonably determine which Gaussians belong to which objects.
@@ 3.2 Object-centric pruning
A key observation in reconstructing a specific object is that the task itself is about a single object. However, directly distilling masks or semantics often leads to the distillation being spread across multiple Gaussians. This is because the Gaussians are typically non-opaque, meaning multiple Gaussians must work together to form surfaces. This, in turn, results in the appearance of floaters that are irrelevant to the object of interest, as shown earlier in Fig. 1. Given that object-specific 3D Gaussians should represent only the target object, we can effectively ‘prune’ these floaters by simply removing those that do not follow the typical connectivity with nearby Gaussians, that is, statistical outliers.
Statistical outlier removal. To do so, we propose a pruning strategy based on statistical outlier removal (Rusu and Cousins 2011). Our core idea is to look into the average distance that a Gaussian forms with its K nearest neighbors, and remove those that do not follow the general trend—those that are in the long tails of the distance distribution. Specifically, denoting the distance between the two Gaussians as $\Delta(\cdot,\cdot)$, we prune a Gaussian $\mathcal{G}_{i}$ if it satisfies:
$\frac{1}{K}\sum_{j\in\mathcal{N}_{i}}\Delta\left(\mathcal{G}_{i},\mathcal{G}_{j}\right)>\bar{\mu}+\alpha\cdot\bar{\sigma},$ (4)
where $\mathcal{N}_{i}$ are the K nearest neighbors, $\bar{\mu}=\frac{1}{NK}\sum_{i}\sum_{j\in\mathcal{N}_{i}}\Delta\left(\mathcal{G}_{i},\mathcal{G}_{j}\right)$ is the average distance and $\bar{\sigma}$ is the standard deviation of distances, and $\alpha$ is a hyperparamter controlling the pruning aggressiveness.
Measuring distance between Gaussians. For Eq. 4 to work properly, it is essential that $\Delta(\cdot,\cdot)$ is defined correctly. A naive implementation would consider only the Gaussian centers, but this quickly leads to poor performance as Gaussians are often elongated to represent edges. We thus propose to use squared 2-Wasserstein distance, $W_{2}^{2}$. Denoting Gaussians as $\mathcal{G}=\mathcal{N}(\mu,\Sigma)$, we thus write:
$\Delta\left(\mathcal{G}_{i},\mathcal{G}_{j}\right)=W_{2}^{2}\!\left(\mathcal{G}_{i},\mathcal{G}_{j}\right)$ (5)
$\;=\lVert\mu_{i}-\mu_{j}\rVert_{2}^{\,2}+\mathrm{Tr}\!\Bigl{(}\Sigma_{i}+\Sigma_{j}-2\bigl{(}\Sigma_{j}^{1/2}\,\Sigma_{i}\,\Sigma_{j}^{1/2}\bigr{)}^{1/2}\Bigr{)},$
where $\Sigma_{j}^{1/2}$ denotes the principal matrix square root of $\Sigma_{j}$; we use the squared form for efficiency.
Implementation. Our CUDA implementation of the pruning process runs within 5 seconds for each scene, which we run once at 500 iterations—the cost is virtually none compared to the time it takes to train Gaussians. This results in an average of 1% of the Gaussians being pruned.